# Labels Matter More Than Models: Quantifying the Benefit of Supervised Time Series Anomaly Detection

### v1

@@ V-A Task Designs
    Analysis of Supervisory Gain: Considering that the previous task might be subject to certain fairness concerns (due to using 50% labeled data), we further reduce the amount of labeled time series used for training (utilizing as little as 10% of the data) and test on the original test set. We use four subsets of the PSM dataset (SP1-SP4) with increasing anomaly counts (from least to most) for this analysis. The specific dataset partitioning method and data characteristics are introduced in Sec. V-B3.
@@ V-B1 Metrics
To assess the models’ event-level anomaly detection capability, we introduce the Aff-F1, UAff-F1 [simad], and Volume Under the Surface for the Precision-Recall surface (VUS-PR) [VUS] scores. Aff-F1 evaluates the interval-level F1-score by utilizing interval membership to quantify the distance between predicted results and ground truth labels. Furthermore, UAff-F1 is an unbiased refinement of Aff-F1, placing greater emphasis on the precision of the model’s detection. VUS-PR is the area under the event-level Precision-Recall curve, which effectively assesses a model’s ability to detect anomalous events.
Furthermore, we introduce Confidence-Consistency Evaluation (CCE) to quantify the prediction consistency of the models [cce]. This metric not only considers the accuracy of the anomaly scores but also accounts for local and global prediction uncertainty. It has been theoretically proven to be more robust and is particularly suitable for scenarios where anomaly labels or anomaly scores are noisy.
@@ V-B2 Baselines
The UTAD-II methods encompass a broader range of techniques, including traditional kernel methods, deep learning models, and Transformer-based approaches, specifically including One-Class Support Vector Machine (OCSVM), Autoencoder (AE), Convolutional Neural Network (CNN), Long Short-Term Memory Network (LSTM), TranAD [trans_ad], USAD [usad], Omni [omni], Anomaly Transformer (A.T.) [anom_trans], TimesNet [timesnet], M2N2 [M2N2], LFTSAD [LFTSAD], and CATCH [catch]. This category integrates shallow kernel-based learning, deep neural networks (for spatial-temporal feature extraction), and state-of-the-art Transformer architectures (for capturing long-range dependencies in time series).
@@ V-B3 Datasets
To evaluate the performance of different algorithms in real-world scenarios, we utilize five multivariate time series anomaly detection datasets: PSM, SWaT, WADI, Swan, and Water [dcdetector, simad]. The characteristics of these datasets are summarized in Table I.
For Task 2, to further investigate the gain conferred by supervision, we select PSM as the base dataset. We partition the original test set’s initial portion to serve as the training set, aiming to verify the impact of supervisory labels on the model. This process results in the four PSM subsets shown in Table II (i.e., SP1, SP2, SP3, and SP4). It is important to note that only PSM satisfies the requirement that anomalies are present throughout the entire test set. The anomaly distribution patterns in the other datasets lack sufficient regularity, making it difficult to partition the dataset—specifically, splitting the entire test set into two continuous segments. Hence, we exclusively chose PSM as the base dataset for this task.
@@ V-D Analysis of Supervisory Gain
To explicitly demonstrate the impact of supervisory information on model performance, we train models using four subsets of the PSM dataset (SP1-SP4) and evaluate them on the original, full PSM test set. Notably, SP1 contains only $10\%$ of the PSM data. The performance of different models on these four subsets is summarized in Table IV.
Strikingly, the average performance of ExtraTrees, when trained on the SP1 subset ($10\%$ of the data), reaches $44.19$, which already surpasses the best UTAD-I method, PCA (average score $44.13$). Moreover, it maintains considerable competitiveness against the best UTAD-II method, CATCH. When $20\%$ of the data (i.e., SP2) is used for training, the average performance of ExtraTrees improves to $53.03$, representing an increase of $14.5\%$ compared to CATCH. Furthermore, when $40\%$ of the data (i.e., SP4) is utilized for training, ExtraTrees shows an average improvement of $62.66\%$ relative to CATCH.
These results provide compelling evidence of the massive gain conferred by supervisory information. They clearly demonstrate that fully exploiting label information yields greater performance benefits than merely pursuing complex model architecture improvements.
TABLE IV: Performance of all supervised methods across the four PSM subsets (SP1–SP4).
Model Source CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR Avg.
PCA PSM 6.56 18.77 57.93 52.38 74.13 55.01 44.13
CATCH PSM 0.62 16.85 84.85 62.92 60.65 51.99 46.31
RF SP1 13.73 15.60 56.82 49.29 66.79 45.75 41.33
SVM 8.28 15.75 45.52 38.47 68.93 47.72 37.45
AdaBoost 6.39 20.37 40.49 36.55 67.03 46.54 36.23
ExtraTrees 17.06 21.56 56.08 47.21 71.68 51.54 44.19
LightGBM 15.63 15.79 63.16 41.29 71.23 50.37 42.91
STAND 17.84 25.16 52.05 35.07 70.46 54.69 42.54
RF SP2 28.75 30.31 25.14 24.83 82.24 64.35 42.60
SVM 15.33 29.60 25.69 21.27 84.38 65.26 40.26
AdaBoost 10.01 18.40 25.37 21.89 76.14 49.52 33.55
ExtraTrees 33.31 30.34 48.76 48.76 85.84 71.13 53.03
LightGBM 32.35 30.38 26.44 25.44 84.74 68.87 44.70
STAND 29.92 27.56 44.93 40.48 84.83 71.37 49.85
RF SP3 33.83 56.66 30.91 29.21 84.85 65.61 50.18
SVM 14.46 4.65 25.62 18.01 83.02 59.35 34.18
AdaBoost 10.78 16.70 27.22 21.90 74.58 48.81 33.33
ExtraTrees 36.45 57.12 53.36 52.30 85.47 66.97 58.61
LightGBM 33.02 22.55 6.82 6.36 81.85 60.27 35.14
STAND 34.44 26.58 42.81 39.48 87.48 69.41 50.03
RF SP4 44.26 63.97 74.74 67.03 90.62 81.21 70.30
SVM 15.66 29.80 22.34 21.79 91.93 80.47 43.66
AdaBoost 11.30 16.11 27.28 22.58 79.42 54.10 35.13
ExtraTrees 50.25 66.67 81.67 76.72 92.56 84.11 75.33
LightGBM 45.30 30.53 12.57 12.57 89.30 78.37 44.77
STAND 42.67 30.46 28.41 28.31 91.14 82.36 50.56
V-E Effectiveness Analysis of STAND
To analyze the effectiveness of the bidirectional hidden state (Bidir.) and the Temporal Encoding Module (TEM) within STAND for anomaly detection, we conducted ablation experiments on these two components. The results are presented in Table V.
When the TEM is removed, the model’s performance significantly degrades across almost all datasets and metrics, with the average score dropping by $37.2\%$. This result underscores that the TEM is the most crucial component of the STAND framework. If the TEM is retained but the Bidir. is disabled, the overall performance decreases by $8.07\%$. This indicates that multi-directional modeling provides additional effective information for the anomaly detection task.
@@ V-E Effectiveness Analysis of STAND
Figure 5: Sensitivity analysis of model performance with respect to three key hyperparameters on four subsets of the PSM dataset. (a)-(b) show the impact of model dimension ($d_{\text{model}}$) on CCE and VUS-PR, respectively; (c)-(d) illustrate the effect of the number of TEM layers; (e)-(f) depict the influence of window size.
V-F Sensitivity Analysis of STAND
@@ V-G Visualization Analysis
Figures 7 and 8 are localized visualizations of Figure 6, depicting two key regions corresponding to the left and right sides of the red dashed line, respectively. The meaning of each subplot within these figures is consistent with that of Figure 6.
@@ VI Conclusion
This paper critically re-evaluates the prevailing research paradigm in time series anomaly detection (TSAD), which heavily favors complex model architectures to address the lack of labels. By proposing STAND, a simple yet effective supervised baseline, we quantify the impact of supervisory signals against architectural sophistication. Our empirical results conclusively show that “labels matter more than models.” Even with a minimal labeling budget (e.g., 10% of data), STAND significantly outperforms existing complex unsupervised methods in terms of detection accuracy, F1 score, and prediction consistency. These findings reveal the limitations of the current algorithm-centric path and highlight the immense value of limited supervision. Consequently, we suggest that future TSAD research should pivot towards data-centric approaches that maximize the utility of available labels to enhance practical deployment performance.
@@ Acknowledgments
[Uncaptioned image] Zhijie Zhong received the B.S. degree in 2022 from the Harbin Engineering University, Harbin, China and he is currently pursuing the Ph.D. degree in the School of Future Technology, South China University of Technology, Guangzhou, China. His research interests include data mining, machine learning, time series analysis, anomaly detection, and large language model (LLM).
[Uncaptioned image] Zhiwen Yu (S’06-M’08-SM’14) received the Ph.D. degree from the City University of Hong Kong, Hong Kong, China, in 2008. He is currently a Professor with the School of Computer Science and Engineering, South China University of Technology, China. He has authored or co-authored more than 200 refereed journal articles and international conference papers, including more than 80 articles in the journals of IEEE Transactions, h-index 52, and Google citation 13800. He is an Associate Editor of IEEE Transactions on Systems, Man, and Cybernetics: Systems. He is a Senior Member of ACM and a member of Council of China Computer Federation (CCF).
[Uncaptioned image] Kaixiang Yang (M’21) received the B.S. degree and M.S. degree from the University of Electronic Science and Technology of China and Harbin Institute of Technology, China, in 2012 and 2015, respectively, and the Ph.D. degree from the School of Computer Science and Engineering, South China University of Technology, China, in 2020. He has been a Research Engineer with the 7th Research Institute, China Electronics Technology Group Corporation, Guangzhou, China, from 2015 to 2017, and has been a Postdoctoral Researcher with Zhejiang University from 2020 to 2021. He is now with the School of Computer Science and Engineering, South China University of Technology. His research interests include pattern recognition, machine learning, and industrial data intelligence.
[Uncaptioned image] C. L. Philip Chen (S’88-M’88-SM’94-F’07) received the M.S. degree in electrical engineering from the University of Michigan, Ann Arbor, MI, USA, in 1985, and the Ph.D. degree in electrical engineering from Purdue University, West Lafayette, IN, USA, in 1988. He is the Chair Professor and the Dean of the School of Computer Science and Engineering, South China University of Technology, Guangzhou, China. Being a Program Evaluator of the Accreditation Board of Engineering and Technology Education (ABET) in USA, for computer engineering, electrical engineering, and software engineering programs, he successfully architects the University of Macau’s Engineering and Computer Science programs receiving accreditations from Washington/Seoul Accord through Hong Kong Institute of Engineers (HKIE), which is considered as his utmost contribution in engineering/computer science education for Macau as the former Dean of the Faculty of Science and Technology. His current research interests include cybernetics, systems, and computational intelligence. Dr. Chen is a fellow of AAAS, IAPR, CAA, and HKIE, and a member of the Academia Europaea (AE), the European Academy of Sciences and Arts (EASA), and the International Academy of Systems and Cybernetics Science (IASCYS). He received IEEE Norbert Wiener Award in 2018 for his contribution in systems and cybernetics, and machine learnings. He is also a highly cited researcher by Clarivate Analytics in 2018 and 2019. He was a recipient of the 2016 Outstanding Electrical and Computer Engineers Award from his alma mater, Purdue University, in 1988. He was the Chair of TC 9.1 Economic and Business Systems of International Federation of Automatic Control from 2015 to 2017 and currently is a Vice President of Chinese Association of Automation (CAA). He was the IEEE Systems, Man, and Cybernetics Society President from 2012 to 2013, the Editor-in-Chief for the IEEE TRANSACTIONS ON SYSTEMS, MAN, AND CYBERNETICS: SYSTEMS from 2014 to 2019, and currently, he is the Editor-in-Chief for the IEEE TRANSACTIONS ON CYBERNETICS, and an Associate Editor of IEEE TRANSACTIONS ON ARTIFICIAL INTELLIGENCE and IEEE TRANSACTIONS ON FUZZY SYSTEMS.
@@
This is the appendix of Labels Matter More Than Models: Quantifying the Benefit of Supervised Time Series Anomaly Detection.

### v2

@@ V-A Task Designs
    Analysis of Supervisory Gain: We further reduce the amount of labeled time series used for training (utilizing as little as 10% of the data) and test on the original test set. We use three subsets of the PSM dataset (SP1-SP3) with increasing anomaly counts (from least to most) for this analysis. The specific dataset partitioning method and data characteristics are introduced in Sec. V-B3.
@@
    Robustness Analysis to Label Noise: To evaluate the reliability of STAD methods under imperfect annotation conditions, we simulate realistic annotator errors by injecting varying proportions of point and segment noise into the training labels, thereby assessing the models’ resilience to corrupted supervisory signals.
  4. 4.
@@ V-B1 Metrics
To assess the models’ event-level anomaly detection capability, we introduce the Aff-F1, UAff-F1 [51], and Volume Under the Surface for the Precision-Recall surface (VUS-PR) [27] scores. Aff-F1 evaluates the interval-level F1-score by utilizing interval membership to quantify the distance between predicted results and ground truth labels. Furthermore, UAff-F1 is an unbiased refinement of Aff-F1, placing greater emphasis on the precision of the model’s detection. VUS-PR is the area under the event-level Precision-Recall curve, which effectively assesses a model’s ability to detect anomalous events.
Furthermore, we introduce Confidence-Consistency Evaluation (CCE) to quantify the prediction consistency of the models [49]. This metric not only considers the accuracy of the anomaly scores but also accounts for local and global prediction uncertainty. It has been theoretically proven to be more robust and is particularly suitable for scenarios where anomaly labels or anomaly scores are noisy.
@@ V-B2 Baselines
The UTAD-II methods encompass a broader range of techniques, including traditional kernel methods, deep learning models, and Transformer-based approaches, specifically including One-Class Support Vector Machine (OCSVM), Autoencoder (AE), Convolutional Neural Network (CNN), Long Short-Term Memory Network (LSTM), TranAD [35], USAD [1], Omni [33], Anomaly Transformer (A.T.) [44], TimesNet [40], M2N2 [16], LFTSAD [6], and CATCH [41]. This category integrates shallow kernel-based learning, deep neural networks (for spatial-temporal feature extraction), and state-of-the-art Transformer architectures (for capturing long-range dependencies in time series).
@@ V-B3 Datasets
To evaluate the performance and scalability of different algorithms in real-world large-scale scenarios, we utilize five well-known multivariate time series anomaly detection datasets collected from complex industrial systems and sensor networks: PSM, SWaT, WADI, Swan, and Water [45, 51]. Notably, datasets like WADI encompass up to 122 highly correlated channels and hundreds of thousands of timestamps, representing typical industrial big data environments. The characteristics of these datasets are summarized in Table I.
For Task 2, to further investigate the gain conferred by supervision, we select PSM as the base dataset. We partition the original test set’s initial portion to serve as the training set. To ensure a fair evaluation and prevent train-test contamination, the evaluation is performed exclusively on the Remaining portion of the data (as denoted in Table II). This process results in the three disjoint PSM subsets shown in Table II (i.e., SP1, SP2, and SP3). It is important to note that only PSM satisfies the requirement that anomalies are present throughout the entire test set. The anomaly distribution patterns in the other datasets lack sufficient regularity, making it difficult to partition the dataset—specifically, splitting the entire test set into two continuous segments. Hence, we exclusively chose PSM as the base dataset for this task.
@@ V-D Analysis of Supervisory Gain
As shown in Table IV, the incorporation of supervisory signals yields substantial performance improvements for STAD models. We can draw several key observations from these results.
First, as the availability of labeled data increases from SP1 to SP3, the performance of supervised methods demonstrates a clear and consistent upward trajectory. Specifically, STAND’s average score improves dramatically from $37.09$ on SP1 to $46.91$ on SP3. Its average rank also ascends from $2.8$ to $1.8$. This highlights the model’s exceptional capability to translate even incremental additions of supervisory information into significant performance gains.
Second, the results reveal a distinct difference among unsupervised methods based on data volume. When the training data is extremely limited (e.g., SP1), traditional UTAD-I methods like PCA actually perform remarkably well. In fact, PCA achieves a highly competitive average rank of $2.2$ on SP1. Conversely, deep learning-based UTAD-II methods, such as CATCH, require massive amounts of data to model normal distributions effectively. As a result, they struggle in low-data regimes. Even as the data volume increases in SP2 and SP3, their overall performance still falls short of the top-tier methods.
Finally, unsupervised methods like CATCH and PCA exhibit stagnant or fluctuating performance across the data subsets. This is because they cannot directly optimize using anomaly labels. In contrast, STAD methods consistently capitalize on the expanded label set. In both SP2 and SP3, STAND achieves the highest average scores ($45.95$ and $46.91$, respectively) and the best average ranks ($1.6$ and $1.8$). This compelling empirical evidence robustly validates our core hypothesis: allocating resources to obtain a small amount of anomaly labels yields far greater performance dividends than relentlessly increasing unsupervised model complexity.
TABLE IV: Performance of all supervised methods across the three PSM subsets (SP1–SP3).
Source Model CCE F1 UAff-F1 AUC-ROC VUS-PR Avg. Score Avg. Rk.
SP1 PCA 5.47 18.17 53.51 73.35 55.78 41.25 2.2
CATCH 0.59 16.89 55.80 62.88 52.47 37.73 4.4
RF 4.64 9.89 34.29 60.37 41.24 30.09 6.8
SVM 6.09 11.69 31.52 63.88 45.37 31.71 5.6
AdaBoost 6.76 15.61 31.95 71.18 49.58 35.02 3.4
ExtraTrees 6.33 16.26 30.89 66.19 47.01 33.34 4.6
LightGBM 4.71 10.15 17.20 66.20 45.81 28.81 6.2
STAND 14.83 19.68 29.74 68.17 53.01 37.09 2.8
SP2 PCA 5.51 18.71 52.88 74.29 57.11 41.70 4.2
CATCH 0.97 17.34 59.94 63.13 54.86 39.25 5.6
RF 10.27 17.00 22.77 71.47 51.47 34.60 6.2
SVM 10.21 17.19 31.41 76.14 55.27 38.04 5.2
AdaBoost 7.85 16.91 22.64 69.73 46.39 32.70 7.4
ExtraTrees 11.19 23.12 39.60 77.26 59.98 42.23 2.4
LightGBM 13.16 21.12 35.69 76.53 57.11 40.72 3.4
STAND 19.76 21.81 42.67 80.48 65.01 45.95 1.6
SP3 PCA 11.09 22.37 51.73 79.03 55.28 43.90 3.8
CATCH 2.60 15.73 53.20 60.61 45.90 35.61 6.4
RF 12.75 15.93 15.97 72.74 50.62 33.60 5.4
SVM 12.25 22.13 52.09 80.42 57.39 44.86 2.8
AdaBoost 10.77 17.95 27.34 71.87 43.73 34.33 6.8
ExtraTrees 11.55 23.60 46.20 78.33 59.07 43.75 3.2
LightGBM 11.43 16.20 35.70 72.06 47.56 36.59 5.8
STAND 24.47 26.54 42.82 80.66 60.09 46.91 1.8
V-E Robustness to Label Noise
In practical industrial environments, acquiring flawlessly accurate anomaly labels is frequently unattainable due to annotator fallibility or inherent system ambiguities. To address potential concerns regarding the reliance of STAD methods on pristine annotations, we empirically investigate the resilience of these models against label noise. We formulate two distinct noise injection paradigms on the PSM dataset to emulate realistic annotation errors:
  1. 1.
    Point Noise: We randomly sample a specified proportion $\rho$ of the labels across the entire training sequence and invert their respective classes (i.e., normal to anomalous, and vice versa). This mechanism simulates sporadic, independent annotation inaccuracies.
  2. 2.
    Segment Noise: We randomly select contiguous segments of normal time steps and reassign their labels to anomalous until the cumulative number of inverted steps reaches the target proportion $\rho$. This paradigm emulates systematic errors, wherein an annotator might misclassify an entire operational phase.
We systematically vary the noise ratio $\rho\in\{0.0,0.01,0.03,0.05,0.07,0.1\}$ and benchmark the STAD models alongside the SOTA UTAD-II method, CATCH. The comparative performance is illustrated in Figure 4.
Figure 4: Performance of supervised models under varying ratios of Point Noise (solid lines) and Segment Noise (dashed lines) on the PSM dataset. The gray dashed line represents the performance of the UTAD-II SOTA method, CATCH, which does not utilize labels and thus remains constant.
Based on the quantitative evaluations, we deduce the following critical observations:
  • Sustained Superiority Under Substantial Noise: Although the performance of all STAD models inherently degrades as the noise ratio escalates, robust methods such as ExtraTrees and STAND preserve a decisive advantage over the SOTA unsupervised baseline (CATCH) across the majority of metrics, even at a severe noise ratio of 10%. For instance, under Segment Noise at $\rho=0.1$, STAND achieves a CCE of $23.47$, an AUC of $80.91$, and a VUS-PR of $61.27$, which still substantially surpass the performance of CATCH ($0.62$, $60.65$, and $51.99$, respectively).
  • Differential Impact of Noise Paradigms: Traditional approaches like Random Forest (RF) manifest pronounced sensitivity to Point Noise, evidenced by a precipitous decline in point-wise metrics such as F1 and Aff-F1. Conversely, Segment Noise exerts a more attenuated degradative impact on these models. Notably, ExtraTrees exhibits remarkable resilience to both noise modalities, consistently maintaining top-tier global performance.
  • Robustness of STAND: The proposed STAND baseline yields stable degradation trajectories, particularly concerning global evaluation metrics (e.g., AUC) and structural consistency metrics (e.g., CCE).
    This empirical evidence suggests that learning a parametric mapping via deep sequence models affords inherent noise tolerance, thereby mitigating severe overfitting to corrupted supervisory signals.
V-F Effectiveness Analysis of STAND
To rigorously validate the architectural design of the STAND framework, we conduct a two-part ablation study: (1) assessing the necessity of the core components, namely the Temporal Encoding Module (TEM) and the bidirectional hidden state (Bidir.), and (2) evaluating the specific choice of the sequential encoder.
V-F1 Component Ablation
Table V presents the ablation results regarding the TEM and Bidir. across five datasets. Removing the TEM effectively reduces the model to a pointwise MLP, resulting in a significant performance degradation across all metrics (e.g., a $37.2\%$ drop in average score). This decline emphasizes that relying solely on pointwise features is insufficient for time-series anomaly detection; the TEM is indispensable for capturing the sequential context that distinguishes anomalous from normal behavior.
Furthermore, retaining the TEM but disabling Bidir. decreases overall performance by $8.07\%$. This indicates that bidirectional modeling yields critical supplementary information. In time-series analysis, anomalies are often characterized not only by preceding events but also by subsequent recovery patterns. Consequently, Bidir. enables STAND to leverage both historical and future contexts, ensuring more robust timestamp-level classification.
@@ V-F1 Component Ablation
V-F2 Impact of Temporal Encoder Architecture
Having established the necessity of a bidirectional temporal encoder, we further investigate the optimality of the BiLSTM architecture within this supervised setting. We compare the default BiLSTM in STAND against two variants: a Bidirectional GRU (w/ GRU) and a Multi-Head Attention mechanism (w/ Attn).
To quantitatively evaluate the performance differences, we conducted a Wilcoxon signed-rank test across the five datasets. As detailed in Table VI, the BiLSTM-based STAND demonstrates statistically significant improvements over both the Attention and GRU variants in terms of the CCE metric ($p<0.05$). Additionally, STAND significantly outperforms the GRU variant on the VUS-PR metric ($p<0.05$).
These statistical findings validate our architectural choice. The inferior performance of the more complex Attention variant suggests that, under conditions of limited supervision, highly parameterized attention mechanisms may be susceptible to overfitting and offer diminishing returns. Conversely, the GRU variant exhibits insufficient representational capacity to consistently localize complex anomalous events, as reflected by the decreased V-PR scores. Ultimately, the BiLSTM architecture provides an optimal balance, delivering stable structural consistency and superior local anomaly detection capabilities without introducing excessive model complexity.
TABLE VI: Performance comparison of different encoder variants across five datasets. Bold indicates the best performance. ‘*‘ and $\dagger$‘ indicate that STAND is statistically significantly better than the w/ Attn and w/ GRU variants, respectively, at the $p<0.05$ level (using the Wilcoxon signed-rank test).
Variant Dataset CCE F1 Aff-F1 UAff-F1 AUC V-PR
w/ Attn PSM 37.46 30.46 17.03 16.96 83.10 71.54
w/ GRU 35.50 30.34 31.80 31.21 85.10 72.63
STAND 46.94 30.33 29.65 29.35 90.74 83.66
w/ Attn SWaT 69.56 58.13 78.21 77.02 92.08 88.08
w/ GRU 52.31 57.94 81.10 80.92 91.29 79.40
STAND 71.48 57.95 87.84 87.42 92.02 87.14
w/ Attn WADI 61.64 77.55 91.17 82.99 93.39 83.22
w/ GRU 60.48 68.07 89.28 82.56 95.20 72.43
STAND 65.24 82.03 89.93 84.07 91.72 87.08
w/ Attn Swan 54.93 24.97 45.68 45.20 87.37 95.58
w/ GRU 50.22 24.76 42.41 41.90 86.07 95.11
STAND 55.53 25.57 46.98 46.65 84.76 97.19
w/ Attn Water 82.28 28.14 98.14 96.29 97.00 24.33
w/ GRU 83.14 32.15 94.90 89.83 98.51 88.71
STAND 84.44 34.01 96.61 93.41 99.62 95.38
w/ Attn Average 61.17 43.85 66.05 63.69 90.59 72.55
w/ GRU 56.33 42.65 67.90 65.28 91.23 81.66
STAND 64.73∗† 45.98 70.20 68.18 91.77 90.09
$p$-value (STAND vs. w/ Attn) 0.043 0.224 0.345 0.224 0.685 0.079
$p$-value (STAND vs. w/ GRU) 0.043 0.079 0.224 0.138 0.892 0.043
@@ V-F2 Impact of Temporal Encoder Architecture
Figure 6: Sensitivity analysis of model performance with respect to three key hyperparameters on three subsets of the PSM dataset. (a)-(b) show the impact of model dimension ($d_{\text{model}}$) on CCE and VUS-PR, respectively; (c)-(d) illustrate the effect of the number of TEM layers; (e)-(f) depict the influence of window size.
V-G Sensitivity Analysis of STAND
@@ V-H Visualization Analysis
Figures 8 and 9 are localized visualizations of Figure 7, depicting two key regions corresponding to the left and right sides of the red dashed line, respectively. The meaning of each subplot within these figures is consistent with that of Figure 7.
@@ VI Conclusion
This paper critically re-evaluates the prevailing research paradigm in time series anomaly detection (TSAD), which heavily favors complex model architectures to address the lack of labels. By proposing STAND, a simple yet effective supervised baseline, we quantify the impact of supervisory signals against architectural sophistication. Our empirical results conclusively show that labels matter more than models. Even with a minimal labeling budget (e.g., 10% of data), STAND significantly outperforms existing complex unsupervised methods in terms of detection accuracy, F1 score, and prediction consistency. These findings reveal the limitations of the current algorithm-centric path and highlight the immense value of limited supervision. Consequently, we suggest that future TSAD research should pivot towards data-centric approaches that maximize the utility of available labels to enhance practical deployment performance. In the future, we plan to extend this data-centric paradigm to distributed big data frameworks and explore active learning strategies to select the most valuable labels from massive data streams efficiently.
@@ Acknowledgments
This is the appendix of Labels Matter More Than Models: Rethinking the Unsupervised Paradigm in Time Series Anomaly Detection.